### v1

# Equality of Hölder exponents for distribution functions of Gibbs measures

Abstract.
We show that the pointwise Hölder spectrum of the distribution function of a Gibbs measure supported on the attractor of a self-conformal subset of the real line satisfies the multifractal formalism.
2020 Mathematics Subject Classification:
37D35, 28A80
1. Introduction
A function $f:\mathbb{R}\to\mathbb{R}$ is said to be Hölder continuous with exponent $\alpha>0$ if there is a constant $C$ such that $|f(x)-f(y)|\leq C|x-y|^{\alpha}$ for all $x$ and $y$. However, this $\alpha$ represents the “worst possible” behavior, and in general, a continuous function can at many points have substantially better regularity than the worst case.
For $\alpha>0$ and $t_{0}\in\mathbb{R}$, write $f\in C^{\alpha}(t_{0})$ if there is a constant $C$ and a polynomial $P$ of degree less than $\alpha$ such that
(1) $|f(t)-P(t-t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t\in\mathbb{R}$}.$
The pointwise Hölder exponent of $f$ at $t$ is the number
(2) $\alpha_{f}(t):=\sup\{\alpha>0:f\in C^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Pointwise Hölder exponents can be difficult to calculate, especially for $\alpha>2$, where, in order to show that $f\not\in C^{\alpha}(t_{0})$, one must prove that no polynomial satisfying (1) exists. For this reason perhaps, many authors use the following, simpler definition of Hölder exponent. Write $f\in\tilde{C}^{\alpha}(t_{0})$ if there is a constant $C$ such that
$|f(t)-f(t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t\in\mathbb{R}$},$
that is, the polynomial $P$ in (1) is constant with value $f(t_{0})$. Define
(3) $\tilde{\alpha}_{f}(t):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Clearly, $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but the reverse inequality may fail in general. For example, if $f(t)=t^{2}$, then $\alpha_{f}(0)=\infty$ (since one can take $P(t)=t^{2}$ in (1) for all $\alpha>2$), but $\tilde{\alpha}_{f}(0)=2$. In addition, a desirable property of Hölder exponents is that they are left unchanged upon perturbation of $f$ by a smooth function $g$. Indeed, $\alpha_{f}(t)=\alpha_{f+g}(t)$, whereas $\tilde{\alpha}_{f}(t)\neq\tilde{\alpha}_{f+g}(t)$ in general.
On the other hand, for certain functions with fractal properties, the two definitions of Hölder exponent coincide. This was shown to be the case by the first author in [1, 2] for certain classes of self-affine functions, including distribution functions of self-similar measures on $\mathbb{R}$ satisfying the open set condition. In this note, we extend this result to distribution functions of arbitrary Gibbs measures on finitely generated self-conformal sets; see the next section for a precise statement.
A function $f$ is multifractal if its pointwise Hölder exponents $\alpha_{f}(t)$ take on more than one value as $t$ ranges through the domain of $f$. For such functions, we are interested in the sets
$E_{f}(\alpha):=\{t\in\mathbb{R}:\alpha_{f}(t)=\alpha\}.$
The Hausdorff and packing multifractal spectra of $f$ are given, respectively, by
$\alpha\mapsto\dim_{H}E_{f}(\alpha)\qquad\mbox{and}\qquad\alpha\mapsto\dim_{P}E_{f}(\alpha),$
where $\dim_{H}$ and $\dim_{P}$ denote Hausdorff and packing dimensions, respectively. If $f=F_{\mu}$ is the distribution function of a Borel probability measure $\mu$, i.e. $F_{\mu}(x):=\mu\left((-\infty,x]\right)$ for $x\in\mathbb{R}$, then by [13, Lemma 5.1] we have
(4) $\tilde{\alpha}_{F_{\mu}}(t_{0})=\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(t_{0},r)\right)}{\log r}.$
For the remainder of this introduction, let $\mu$ a Gibbs measure associated with an arbitrary Hölder potential on the attractor of a conformal iterated function system (IFS) satisfying the open set condition - see the next section for precise definitions. It is well known that the sets
$\mathscr{E}(\alpha;\mu):=\left\{x\in\mathbb{R}\Big{|}\lim_{r\rightarrow 0}\frac{\log\mu\left(B(x,r)\right)}{\log r}=\alpha\right\},\qquad\alpha\in\mathbb{R},$
satisfy the multifractal formalism; that is, $\dim_{H}\mathscr{E}(\alpha;\mu)=\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$, where $\beta^{*}(\alpha)$ is the Legendre transform of a suitable scaling function $\beta(q)$, defined in the next section. The second author and Sumi [13] showed recently that, in the case when all of the maps in the IFS are strictly increasing, the Hausdorff spectrum of the modified level sets
$\mathscr{E}_{*}(\alpha;\mu):=\left\{x\in\mathbb{R}\Big{|}\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(x,r)\right)}{\log r}=\alpha\right\},\qquad\alpha\in\mathbb{R},$
is also given by the multifractal formalism: $\dim_{H}\mathscr{E}_{*}(\alpha;\mu)=\beta^{*}(\alpha)$. It is straightforward to extend this result to general conformal IFS. Thus, as a consequence of our main theorem, the pointwise Hölder spectrum of $F_{\mu}$ is given by
$\dim_{H}E_{F_{\mu}}(\alpha)=\beta^{*}(\alpha).$
In a forthcoming paper, we will show that the situation is different for the packing spectrum.
2. Notation and definitions
Let $X\subset\mathbb{R}$ be a compact interval. A $\mathcal{C}^{1+\epsilon}$ conformal iterated function system on $X$ is given by an index set $I$ with $2\leq\#I<\infty$ and $\mathcal{C}^{1+\epsilon}$ conformal contracting maps $\Phi=(\phi_{i})_{i\in I}$, $\phi_{i}:X\rightarrow X$.
More precisely, we mean that each $\phi_{i}$ extends to a $\mathcal{C}^{1+\epsilon}$ differentiable map on a neighborhood of $X$ which satisfies $0<\left|\phi_{i}^{\prime}(x)\right|<1$ for every $x\in X$ and for each $i\in I$. For ease of presentation, we assume that each of the maps $\phi_{i},i\in I$ is strictly increasing. However, this restriction is not essential. We refer to [18] for further details and basic properties of conformal iterated function systems. Let $\pi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ denote the coding map of $\Phi$ which is for $\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}$ given by
$\bigcap_{n\geq 1}\phi_{\omega_{1}}\circ\dots\circ\phi_{\omega_{n}}(X)=\left\{\pi(\omega)\right\}.$
Let $\Lambda:=\pi(I^{\mathbb{N}})$; then $\Lambda$ satisfies the set equation $\Lambda=\bigcup_{i\in I}\phi_{i}(\Lambda)$. We refer to $\Lambda$ as the self-conformal set generated by $\Phi$. We say that $\Phi$ satisfies the open set condition if $\phi_{i}(\mathrm{Int}(X))\cap\phi_{j}(\mathrm{Int}(X))=\varnothing$ for all $i,j\in I$ with $i\neq j$. To utilize the symbolic thermodynamic formalism (see [6]) we will also need the following definitions. We denote by $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ the left-shift on $I^{\mathbb{N}}$ which becomes a compact metric space endowed with the shift metric. Let $\varphi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be the geometric potential of $\Phi$, given by
(5) $\varphi(\omega):=\log\left(\phi_{\omega_{1}}^{\prime}(\pi(\sigma(\omega)))\right),\quad\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}.$
Since each $\phi_{i}$ is $\mathcal{C}^{1+\epsilon}$ differentiable, it follows that $\varphi$ is Hölder continuous. Let $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be Hölder continuous and denote by
$S_{n}\psi:=\sum_{k=0}^{n-1}\psi\circ\sigma^{k},\qquad n\geq 1$
its ergodic sums. We say that a Borel probability measure $\mu_{\psi}$ on $I^{\mathbb{N}}$ is a Gibbs measure for $\psi$ if there exist two constants $C\geq 1$ and $P\in\mathbb{R}$ such that for each $n\in\mathbb{N}$ and for all $w\in I^{n}$ we have
(6) $C^{-1}\leq\frac{\mu_{\psi}[w_{1},\dots,w_{n}]}{e^{S_{n}\psi(\overline{w})-nP}}\leq C.$
Here $\overline{w}$ denotes the infinite periodic sequence with period block $w$. It is well known that a Gibbs measure $\mu_{\psi}$ for a Hölder continuous potential always exists [6]. Further note that the constant $P$ in the definition of a Gibbs measure is equal to the topological pressure of $\psi$ with respect to $\sigma$, defined by
$\mathcal{P}(\psi):=\lim_{k\to\infty}\frac{1}{k}\log\sum_{w\in I^{k}}\exp S_{k}\psi(\overline{w}).$
We will frequently make use of the fact that any Hölder continuous $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ satisfies the bounded distortion property
(7) $\sup_{n\in\mathbb{N}}\sup_{w\in I^{n}}\sup_{x,y\in I^{\mathbb{N}}}\left|S_{n}\psi(w\rho)-S_{n}\psi(w\tau)\right|<+\infty.$
We remark that if $\mu_{\psi}$ is a Gibbs measure for $\psi$, then it is also a Gibbs measure for $\psi-\mathcal{P}(\psi)$. We may therefore always assume that $\mathcal{P}(\psi)=0$ by replacing $\psi$ with $\psi-\mathcal{P}(\psi)$.
Let
(8) $\alpha_{+}:=\sup_{\omega\in I^{\mathbb{N}}}\limsup_{n\rightarrow\infty}\frac{S_{n}\psi(\omega)}{S_{n}\varphi(\omega)},\qquad\alpha_{-}:=\inf_{\omega\in I^{\mathbb{N}}}\liminf_{n\rightarrow\infty}\frac{S_{n}\psi(\omega)}{S_{n}\varphi(\omega)}.$
It is well known that the multifractal spectrum $\mathcal{F}(\alpha)$ is complete [28], that is, $\mathcal{F}(\alpha)\neq\varnothing$ if and only if $\alpha\in\left[\alpha_{-},\alpha_{+}\right]$. Moreover, the following dichotomy is well known [25]: either we have $\alpha_{-}<\alpha_{+}$, or $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$. Note that in the latter case we have $\alpha_{-}=\alpha_{+}=\dim_{H}\Lambda$.
Since $\varphi<0$, there is for each $q\in\mathbb{R}$ a unique $\beta(q)$ such that
(9) $\mathcal{P}(\beta(q)\varphi+q\psi)=0.$
We denote by
(10) $\beta^{*}(\alpha):=\inf_{q\in\mathbb{R}}\left\{\beta(q)+\alpha q\right\}$
the Legendre transform of $\beta$. If $\alpha_{-}<\alpha_{+}$, then it is well known that $\beta^{*}$ is real-analytic and strictly concave on $(\alpha_{-},\alpha_{+})$ [23].
3. The main theorem
Our main result is the following:
Theorem 3.1.
Let $\Phi=(\phi_{i}:X\rightarrow X)_{i\in I}$ be a conformal iterated function system satisfying the open set condition. Let $\varphi$ denote the geometric potential of $\Phi$ defined in (5) and let $\psi$ be any Hölder potential with $\mathcal{P}(\psi)=0$ such that $\psi$ is not cohomologous to $\varphi$. Let $\mu_{\psi}$ be a Gibbs measure for $\psi$. Then, for the distribution function $F$ of $\mu_{\psi}\circ\pi^{-1}$, we have that $\alpha_{F}=\tilde{\alpha}_{F}$.
Remark 3.2.
The condition that $\psi$ is not cohomologous to $\varphi$ is needed to rule out some degenerate cases. As a simple example, consider the IFS $\Phi=\{x/2,(x+1)/2\}$, whose attractor is the interval $[0,1]$. Clearly, $\varphi\equiv-\log 2$. Choosing $\psi\equiv-\log 2$ as well, we see that $\mu:=\mu_{\psi}\circ\pi^{-1}$ is Lebesgue measure on $[0,1]$, so $F$ is infinitely differentiable on $(0,1)$ and hence $\alpha_{F}(t)\equiv\infty$ on $(0,1)$. On the other hand, it is easy to see that $\tilde{\alpha}_{F}(t)\equiv 1$ on $(0,1)$.
More generally, we claim that for a Hölder continuous $\psi$ with $\mathcal{P}(\psi)=0$ we have that $\psi$ is cohomologous to $\varphi$ if and only if $\Lambda$ is an interval and some/any Gibbs measure $\mu_{\psi}$ for $\psi$ satisfies that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to the Lebesgue measure with a Radon-Nykodym density bounded away from zero and infinity. To prove this, first assume that $\psi$ is cohomologous to $\varphi$. Hence, $P(\varphi)=P(\psi)=0$ and thus, $\dim_{H}(\Lambda)=1$. Hence, by the well-known construction of conformal measures (see e.g. [19, Theorem 4.2.11]), $\Lambda$ has positive Lebesgue measure. Hence, by Lebesgue’s density theorem, we conclude that $X=\bigcup_{i\in I}\Phi_{i}(X)$ and thus, $\Lambda=X$ is an interval. Indeed, if $X\setminus\bigcup_{i\in I}\Phi_{i}(X)$ was non-empty, it would contain an interval of positive length. By standard arguments, this would imply that $\Lambda$ contains no Lebesgue density points, which gives the desired contradiction. Consequently, the normalised Lebesgues measure on $\Lambda$ is a Gibbs measure for $\varphi$. Since $\psi$ is cohomologous to $\varphi$, $\mu_{\psi}$ is also a Gibbs measure for $\varphi$. It follows that $\mu_{\psi}$ is equivalent to the Lebesgue measure with a density bounded away from zero and infinity.
Conversely, assume that $\Lambda$ is an interval and $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to Lebesgue with density bounded away from zero and infinity. Hence, Lebesgue measure is also Gibbs measure for $\psi$. Since Lebesgue measure is also a Gibbs measure for $\varphi$ and $P(\varphi)=P(\psi)=0$, it follows from [19, Theorem 2.2.7] that $\varphi$ is cohomologous to $\psi$.
The key to proving Theorem 3.1 is to prove Proposition 3.4 below. The proof of that proposition uses the following lemma.
Lemma 3.3.
Suppose that $P(\psi)=0$. Suppose that there exists a finite set $A\subset\mathbb{R}$ and $\ell_{0}\geq 1$ such that for all $\ell\geq\ell_{0}$ and all $k_{1},\dots,k_{\ell}\in I$ we have
$\frac{S_{\ell}\psi\left(\overline{k_{1}\dots k_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{k_{1}\dots k_{\ell}}\right)}\in A.$
Then $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $A=\left\{\dim_{H}\Lambda\right\}$.
Proof.
By the bounded distortion property (12), it is easy to see that for each $\ell_{0}\geq 1$,
$\alpha_{-},\alpha_{+}]=\overline{\left\{\frac{S_{\ell}\psi\left(\overline{k_{1}\dots k_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{k_{1}\dots k_{\ell}}\right)}:\ell\geq\ell_{0},k_{1},\dots,k_{\ell}\in I\right\}}.$
Proposition 3.4.
Assume that $\mathcal{P}(\psi)=0$, and $\psi$ is not cohomologous to $\varphi$. Then, for every $k\in\mathbb{N}$ and $x\in\mathbb{R}$ the limit
$\lim_{y\rightarrow x}\frac{F(y)-F(x)}{(y-x)^{k}}$
either does not exist, or it takes the value 0 or $\infty$.
Proof of Proposition 3.4.
The statement is trivial for even $k$, since $F$ is increasing. So below we will assume that $k$ is odd.
Suppose by way of contradiction that
$\lim_{y\rightarrow x}\frac{F(y)-F(x)}{(y-x)^{k}}=L\in(0,+\infty).$
Note that the limit cannot be negative, since $F$ is increasing.
Let $\omega\in I^{\mathbb{N}}$ such that $\pi(\omega)=x$. Let $s_{n},t_{n}]:=\pi\left(\left[\omega_{n}\right]\right)$. Then $s_{n}\leq x\leq t_{n}$, and
(11) $\begin{split}\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}&=\left(\frac{t_{n}-x}{t_{n}-s_{n}}\right)^{k}\cdot\frac{F(t_{n})-F(x)}{(t_{n}-x)^{k}}\\ &\quad+\left(\frac{x-s_{n}}{t_{n}-s_{n}}\right)^{k}\cdot\frac{F(s_{n})-F(x)}{(s_{n}-x)^{k}}.\end{split}$
Denote
$r_{n}:=\frac{t_{n}-x}{t_{n}-s_{n}}.$
Observe that $r_{n}\in[0,1]$ and
$1-r_{n}=\frac{x-s_{n}}{t_{n}-s_{n}}.$
There exists $\ell\in\mathbb{N}$ and $\tau_{1}\dots\tau_{\ell}\in I^{\ell}$ containing at least two different letters such that
$S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)\neq 0.$
Indeed such $\tau_{1}\dots\tau_{\ell}$ always exist for otherwise, Lemma 3.3 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=k=1$, contradicting our hypothesis.
Taking a subsequence if necessary, we may assume that $\lim r_{n}=r\in[0,1]$. Then by (11), we have
$\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}\rightarrow\left(r^{k}+(1-r)^{k}\right)L\in\left[2^{1-k}L,L\right],$
as routine calculus shows that the function $r\mapsto r^{k}+(1-r)^{k}$ is minimized at $r=1/2$. Hence,
(12) $\lim_{n\rightarrow\infty}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)^{k}}=\lim_{n\rightarrow\infty}\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}\in\left[2^{1-k}L,L\right].$
Now, we fix an integer $N\geq 2$ and consider the interval
$\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)=:\left[s_{n}^{N},t_{n}^{N}\right].$
Put
$r_{n}^{N}:=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}.$
Recall that the maps $\phi_{i},i\in I$ are all increasing, so if $u$ and $v$ are two words of the same length and $u\prec v$ in the lexicographical order, then $\pi([u])$ lies to the left of $\pi([v])$.
We claim that there is an infinite subset $\mathcal{N}\subset\mathbb{N}$ such that for each $n\in\mathcal{N}$, the point $x$ and the interval $s_{n}^{N},t_{n}^{N}]$ are separated by at least a full cylinder interval of level at most $n+\ell+1$. To see this, let $j<\ell$ be the smallest integer such that $\tau_{j+1}\neq\tau_{1}$. Such $j$ exists since $\tau_{1}\dots\tau_{\ell}$ contains at least two different letters. Write $\tau_{1}\dots\tau_{\ell}=\tau_{1}^{j}\tau_{j+1}\dots\tau_{\ell}$, and consider two cases:
Case 1. If $\omega$ ends in $\tau_{1}^{\infty}$, then for every sufficiently large $n\in\mathbb{N}$ the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}^{j+1}\tau_{j+1}]\big{)}$ separates $x$ from the interval $\pi\big{(}[\omega|_{n}\tau_{1}\dots\tau_{j+1}]\big{)}$, and hence from $s_{n}^{N},t_{n}^{N}]$.
Case 2. Otherwise, there are infinitely many $n$ such that $\omega_{n+1}\neq\tau_{1}$. Take such an $n$. If $\omega_{n+1}<\tau_{1}$, then
$\omega_{n+1}\dots\omega_{n+\ell+1}\prec\tau_{1}1^{\ell}\prec\tau_{1}\dots\tau_{\ell}\tau_{1},$
so the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}1^{\ell}]\big{)}$ lies between $x$ and $s_{n}^{N},t_{n}^{N}]$. A similar argument, with the inequalities reversed, applies when $\omega_{n+1}>\tau_{1}$.
In either case, the claim follows. By replacing $\mathcal{N}$ by an infinite subset if necessary, we may assume that either $t_{n}^{N}>x$ for all $n\in\mathcal{N}$, or $t_{n}^{N}<x$ for all $n\in\mathcal{N}$. Without loss of generality, we assume the former. The argument in the latter case is similar. Thus, $r_{n}^{N}>0$ for all $n\in\mathcal{N}$.
Let $r_{\min}>0$ be a number such that $|\phi_{i}^{\prime}(x)|\geq r_{\min}$ for all $i\in I$ and $x\in X$. Then
(13) $\mathrm{diam}\pi([uv])\geq\mathrm{diam}\pi([u])\cdot r_{\min}^{|v|}$
for any two finite words $u$ and $v$.
For $n\in\mathcal{N}$, let $u_{n},v_{n}]$ be a cylinder interval of length $n+\ell+1$ separating $x$ and $s_{n}^{N},t_{n}^{N}]$. Then by (13),
$|t_{n}^{N}-x|\geq v_{n}-u_{n}\geq\mathrm{diam}\pi([\omega|_{n}])\cdot r_{\min}^{\ell+1}.$
Hence, there is a constant $C=C(\ell)>1$, depending only on $\ell$ but not on $n$ or $N$, such that
$C^{-1}<\frac{t_{n}^{N}-x}{\mathrm{diam}\big{(}\pi\left(\left[\omega|_{n}\right]\right)\big{)}}<C,$
for all $n\in\mathcal{N}$. We denote this by
(14) $t_{n}^{N}-x\asymp_{\ell}\mathrm{diam}\big{(}\pi\left(\left[\omega|_{n}\right]\right)\big{)}.$
Now (14) implies
(15) $r_{n}^{N}=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}\asymp_{\ell}\frac{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)}{\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}\asymp_{\ell}e^{-NS_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
Hence, by passing to a subsequence, using a diagonal argument, we may assume that $r_{n}^{N}$ converges to some number $r^{N}\in\mathbb{R}$ for every $N\geq 1$.
Using (11) and letting $n\rightarrow\infty$ we obtain
(16) $\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}=\left((r^{N})^{k}+(1-r^{N})^{k}\right)L\geq 2^{1-k}L.$
On the other hand, by the Gibbs property (6) of $\mu_{\psi}$ and the bounded distortion property (12) we have
$\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}$ $=\frac{\mu_{\psi}\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)\right)^{k}}$
$\asymp_{\ell}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}\right]\right)\right)^{k}}\cdot e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}$
$\asymp_{\ell}e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
If $k=1$ this already gives a contradiction, since we then have equality in (16), which implies $S_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=0$.
Suppose now that $k\geq 3$. We have shown that
$2^{1-k}L\leq\big{(}(r^{N})^{k}+(1-r^{N})^{k}\big{)}L=\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}\asymp e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
Since $S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)\neq 0$ it follows that $S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)>0$, and hence, $r^{N}\rightarrow\infty$ as $N\rightarrow\infty$. Since $k$ is odd, $r^{k}+(1-r)^{k}=kr^{k-1}+O(r^{k-2})$ as $r\to\infty$, so we conclude that $(r^{N})^{k}+(1-r^{N})^{k}\asymp(r^{N})^{k-1}$. Thus,
$(r^{N})^{k-1}\asymp e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
Together with (15), this implies
$e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}\asymp(r^{N})^{k-1}\asymp e^{-N(k-1)S_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})},$
and hence,
$S_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=-(k-1)S_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}}),$
i.e.
$S_{\ell}\left(\psi-\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=0.$
We have thus shown that if $\tau_{1}\dots\tau_{\ell}$ contains at least two different letters, then
$S_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})\neq 0\implies S_{\ell}\left(\psi-\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})=0.$
Therefore,
(17) $\frac{S_{\ell}\psi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)}\in\left\{1,k\right\}.$
It remains to consider the fixed points $\overline{\tau_{1}}$:
$\frac{S_{\ell}\psi\left(\overline{\tau_{1}}\right)}{S_{\ell}\varphi\left(\overline{\tau_{1}}\right)}=\frac{\psi\left(\overline{\tau_{1}}\right)}{\varphi\left(\overline{\tau_{1}}\right)}.$
Hence, the set
$A:=\left\{\frac{S_{\ell}\psi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)},\quad\ell\geq 1,\quad\tau_{1}\dots\tau_{\ell}\in I\right\}$
contains only finitely many values, at least one of which is 1 or $k$ in view of (17). Now, Lemma 3.3 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=1$. Hence, $\varphi$ is cohomologous to $\psi$, which is a contradiction. ∎
Proof of Theorem 3.1.
We show that, for any $\alpha>1$, if there is a polynomial $P$ of degree $n<\alpha$ such that
(18) $|F(t)-P(t-t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t$},$
then $P$ is constant: $P(t)=f(t_{0})$ for every $t\in\mathbb{R}$. This clearly implies that $\alpha_{F}(t_{0})=\tilde{\alpha}_{F}(t_{0})$.
Suppose, to the contrary, that $P$ has degree $1\leq n<\alpha$ and satisfies (18). We can write
$P(t)=a_{0}+a_{1}t+a_{2}t^{2}+\dots+a_{n}t^{n},$
where $a_{0}=P(0)=F(t_{0})$. We will show by induction that $a_{1}=a_{2}=\dots=a_{n}=0$. First, write
$\frac{F(t)-F(t_{0})}{t-t_{0}}=\frac{F(t)-P(t-t_{0})}{t-t_{0}}+\frac{P(t-t_{0})-F(t_{0})}{t-t_{0}},$
and observe by (18) that
$\left|\frac{F(t)-P(t-t_{0})}{t-t_{0}}\right|\leq C|t-t_{0}|^{\alpha-1}\to 0\qquad\mbox{as $t\to t_{0}$}.$
Hence,
$F^{\prime}(t_{0})=\lim_{t\to t_{0}}\frac{F(t)-F(t_{0})}{t-t_{0}}=\lim_{t\to t_{0}}\frac{P(t-t_{0})-P(0)}{t-t_{0}}=P^{\prime}(0)=a_{1}.$
By Proposition 3.4, it follows that $a_{1}=0$, establishing the basis for the induction.
Next, take $2\leq j\leq n$, and assume we have already shown that $a_{1}=\dots=a_{j-1}=0$. Thus,
$P(t)=F(t_{0})+a_{j}t^{j}+\dots+a_{n}t^{n}.$
Since $j\leq n<\alpha$, we can apply the same idea as above to conclude that
$\lim_{t\to t_{0}}\frac{F(t)-F(t_{0})}{(t-t_{0})^{j}}=\lim_{t\to t_{0}}\frac{P(t-t_{0})-P(0)}{(t-t_{0})^{j}}=\lim_{t\to t_{0}}\sum_{i=j}^{n}\frac{a_{i}(t-t_{0})^{i}}{(t-t_{0})^{j}}=a_{j},$
and therefore, $a_{j}=0$ by Proposition 3.4.
By induction, we have established the claim, so $P(t)\equiv f(t_{0})$, as desired. ∎
4. The pointwise Hölder spectrum of $F_{\mu}$
Recall from the Introduction that $E_{F}(\alpha)=\{t\in\mathbb{R}:\alpha_{F}(t)=\alpha\}$. As a consequence of Theorem 3.1, we obtain the following result.
Corollary 4.1.
With the notation of Theorem 3.1, we have that
$\dim_{H}E_{F}(\alpha)=\beta^{*}(\alpha)\qquad\forall\,\alpha\in[\alpha_{-},\alpha_{+}],$
whereas $E_{F}(\alpha)=\varnothing$ for $\alpha\not\in[\alpha_{-},\alpha_{+}]$.
Proof.
By (4) and Theorem 3.1, it follows that
$E_{F}(\alpha)=\mathscr{E}_{*}(\alpha,\mu_{\psi}).$
As shown in [13], $\dim_{H}\mathscr{E}_{*}(\alpha,\mu_{\psi})=\beta^{*}(\alpha)$ for $\alpha\in[\alpha_{-},\alpha_{+}]$, and $\mathscr{E}_{*}(\alpha,\mu_{\psi})=\varnothing$ otherwise. ∎
Remark 4.2.
The situation is different for the packing spectrum. We illustrate this for the case when $\mu$ is a self-similar measure. Suppose the maps $\{\phi_{i}\}_{i\in I}$ are similarities with contraction ratios $\{r_{i}\}$, and $\mathbf{p}=(p_{1},\dots,p_{m})$ is a probability vector. Put $\psi(\omega)=\log p_{\omega_{1}}$ for $\omega\in I^{\mathbb{N}}$. Likewise we obtain $\varphi(\omega)=\log r_{\omega_{1}}$. Then $\mu_{\psi}$ is just the self-similar measure satisfying
$\mu=\sum_{i\in I}p_{i}\mu\circ\phi_{i}^{-1},$
and the pressure equation (9) becomes
$\sum_{i\in I}r_{i}^{\beta(q)}p_{i}^{q}=1.$
Let $A(x;\mu)$ denote the set of accumulation points of $\log\mu(B(x,r))/\log r$ as $r\to 0$. It was shown by Baek, Olsen and Snigireva [3] under the strong separation condition, and later by Zhou and Chen [30] under the open set condition, that for any closed interval $I$, the set
$\Delta(I):=\{x\in\mathbb{R}:A(x;\mu)=I\}$
has packing dimension
$\dim_{P}\Delta(I)=\max_{\alpha\in I}\beta^{*}(\alpha).$
Let $\alpha_{0}$ be the unique value of $\alpha$ maximizing $\beta^{*}(\alpha)$. Take first $\alpha<\alpha_{0}$, and put $I=[\alpha,\alpha_{0}]$. Since $\mathscr{E}_{*}(\alpha)\supset I$, it follows that
$\dim_{P}E_{F}(\alpha)=\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\geq\dim_{P}\Delta(I)=\beta^{*}(\alpha_{0})=\dim_{P}\Lambda.$
The reverse inequality is trivial since $E_{F}(\alpha)\subset\Lambda$. Hence, $\dim_{P}E_{F}(\alpha)=\beta^{*}(\alpha_{0})$.
For $\alpha\geq\alpha_{0}$, we trivially have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\geq\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$ since $\mathscr{E}_{*}(\alpha;\mu)\supset\mathscr{E}(\alpha,\mu)$. But we also have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\leq\beta^{*}(\alpha)$. This follows directly from a careful examination of the proof that $\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$; see, for instance, [21, p. 501]. We thus conclude that
(19) $\dim_{P}E_{F}(\alpha)=\begin{cases}\beta^{*}(\alpha_{0})&\mbox{if $\alpha_{-}\leq\alpha\leq\alpha_{0}$},\\ \beta^{*}(\alpha)&\mbox{if $\alpha_{0}\leq\alpha\leq\alpha_{+}$}.\end{cases}$
Note that, on the left half of the multifractal spectrum, the Hausdorff dimension does not differentiate between the sets $\mathscr{E}(\alpha;\mu)$ and $\mathscr{E}_{*}(\alpha;\mu)$, but the packing dimension does. Also, it is unusual to see a packing spectrum that is constant on one side of the maximum.
In a forthcoming paper, we will extend the packing spectrum (19) to the case of distribution functions of general Gibbs measures.

### v2

@@
 Abstract.
-We show that the pointwise Hölder spectrum of the distribution function of a Gibbs measure supported on the attractor of a self-conformal subset of the real line satisfies the multifractal formalism.
+Pointwise Hölder exponents describe the degree of regularity of a function near a point.
+For a function $f:\mathbb{R}\to\mathbb{R}$, a number $\alpha>0$ and a point $t_{0}\in\mathbb{R}$, write $f\in C^{\alpha}(t_{0})$ if there is a constant $C$ and a polynomial $P$ of degree less than $\alpha$ such that
+$|f(t)-P(t-t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t\in\mathbb{R}$}.$
+The pointwise Hölder exponent of $f$ at $t_{0}$ is the number
+$\alpha_{f}(t_{0}):=\sup\{\alpha>0:f\in C^{\alpha}(t_{0})\}.$
+A simpler quantity, also frequently called pointwise Hölder exponent in the mathematical literature, is the number
+$\tilde{\alpha}_{f}(t_{0}):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t_{0})\},$
+where $f\in\tilde{C}^{\alpha}(t_{0})$ means that there is a constant $C>0$ such that $|f(t)-f(t_{0})|\leq C|t-t_{0}|^{\alpha}$ for all $t\in\mathbb{R}$.
+Clearly $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but strict inequality is possible and in fact common.
+In this paper we consider the case when $f=F_{\mu}$ is the distribution function of a Gibbs measure $\mu$ associated with an arbitrary Hölder continuous potential $\psi$ on a self-conformal set, and show that, under a very mild condition on $\psi$, $\alpha_{f}(t)=\tilde{\alpha}_{f}(t)$ for all $t$.
+As a consequence, we deduce that the pointwise Hölder spectrum of $f$ satisfies the multifractal formalism.
+As an application, we derive the pointwise Hölder spectrum of conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ maps of an interval.
+Key words and phrases:
+Hölder exponent, Gibbs measure, Conformal iterated function system, Distribution function, Multifractal spectrum, Bounded distortion property
 2020 Mathematics Subject Classification:
-37D35, 28A80
+Primary: 26A30, 37D35. Secondary: 28A80
 1. Introduction
@@
 The pointwise Hölder exponent of $f$ at $t$ is the number
-(2) $\alpha_{f}(t):=\sup\{\alpha>0:f\in C^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
+$\alpha_{f}(t):=\sup\{\alpha>0:f\in C^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
+Of great interest are the level sets
+$E_{f}(\alpha):=\{t\in\mathbb{R}:\alpha_{f}(t)=\alpha\},\qquad\alpha\geq 0.$
+The multifractal spectrum, or pointwise Hölder spectrum, of $f$ is the function
+$\alpha\mapsto\dim_{H}E_{f}(\alpha),$
+where $\dim_{H}$ denotes Hausdorff dimension.
+Hölder spectra are an important analytical tool in the study of certain physical processes that exhibit a wide range of local regularity behavior.
+They were introduced by Mandelbrot [15] and Frisch and Parisi [8] to study intermittent turbulence flows and intensity of seismic waves.
+Jaffard [10, 11] determined Hölder spectra for a class of self-similar functions satisfying a certain smoothness condition from a mathematical point of view using wavelet theory.
+Subsequent authors [4, 5, 12, 23] computed Hölder spectra of specific self-affine functions; their results were generalized by the first author in [1, 2].
+For pointwise Hölder spectra of Weierstrass-type functions, we refer to [18].
 Pointwise Hölder exponents can be difficult to calculate, especially for $\alpha>2$, where, in order to show that $f\not\in C^{\alpha}(t_{0})$, one must prove that no polynomial satisfying (1) exists.
@@
 that is, the polynomial $P$ in (1) is constant with value $f(t_{0})$. Define
-(3) $\tilde{\alpha}_{f}(t):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
+$\tilde{\alpha}_{f}(t):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
 Clearly, $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but the reverse inequality may fail in general.
-For example, if $f(t)=t^{2}$, then $\alpha_{f}(0)=\infty$ (since one can take $P(t)=t^{2}$ in (1) for all $\alpha>2$), but $\tilde{\alpha}_{f}(0)=2$.
+For example, if $f(t)=t^{2}$, then $\alpha_{f}(t_{0})=\infty$ for all $t_{0}\in\mathbb{R}$ (since one can take $P(t)=(t+t_{0})^{2}$ in (1) for all $\alpha>2$), but $\tilde{\alpha}_{f}(0)=2$ whereas $\tilde{\alpha}_{f}(t_{0})=1$ for all $t_{0}\neq 0$.
 In addition, a desirable property of Hölder exponents is that they are left unchanged upon perturbation of $f$ by a smooth function $g$.
 Indeed, $\alpha_{f}(t)=\alpha_{f+g}(t)$, whereas $\tilde{\alpha}_{f}(t)\neq\tilde{\alpha}_{f+g}(t)$ in general.
-On the other hand, for certain functions with fractal properties, the two definitions of Hölder exponent coincide.
-This was shown to be the case by the first author in [1, 2] for certain classes of self-affine functions, including distribution functions of self-similar measures on $\mathbb{R}$ satisfying the open set condition.
-In this note, we extend this result to distribution functions of arbitrary Gibbs measures on finitely generated self-conformal sets; see the next section for a precise statement.
-A function $f$ is multifractal if its pointwise Hölder exponents $\alpha_{f}(t)$ take on more than one value as $t$ ranges through the domain of $f$. For such functions, we are interested in the sets
-$E_{f}(\alpha):=\{t\in\mathbb{R}:\alpha_{f}(t)=\alpha\}.$
-The Hausdorff and packing multifractal spectra of $f$ are given, respectively, by
-$\alpha\mapsto\dim_{H}E_{f}(\alpha)\qquad\mbox{and}\qquad\alpha\mapsto\dim_{P}E_{f}(\alpha),$
-where $\dim_{H}$ and $\dim_{P}$ denote Hausdorff and packing dimensions, respectively.
+Unfortunately, the two definitions of pointwise Hölder exponent have frequently been conflated in the literature, often without proper justification.
+For this reason we feel that it is critically important to study classes of functions, ideally as general as possible, for which the two notions of Hölder exponent coincide.
+In general, if $f\in C^{\alpha}(t_{0})$ for $\alpha>1$ and $f^{\prime}(t_{0})\neq 0$, then $\tilde{\alpha}_{f}(t_{0})=1<\alpha_{f}(t_{0})$.
+Turning this around, it follows that if $\alpha_{f}(t_{0})=\tilde{\alpha}_{f}(t_{0})$, then either $\alpha_{f}(t_{0})\leq 1$ or $f^{\prime}(t_{0})=0$.
+This shows that the two definitions of Hölder exponent disagree for most functions known from calculus.
+On the other hand, many sufficiently irregular functions have the property that the only possible finite value of their derivative is zero.
+For example, it was shown by the first author in [1, 2] that for certain classes of self-affine functions, including distribution functions of self-similar measures on $\mathbb{R}$ satisfying the open set condition, the two definitions of pointwise Hölder exponent coincide.
+In this note, we extend this result to distribution functions of arbitrary Gibbs measures on finitely generated self-conformal sets; see Theorem 3.1 for a precise statement.
+Although it may seem plausible that the result can be extended from self-similar measures to self-conformal ones, the proof nonetheless requires an entirely new approach.
+This is because the method of proof in [2], using divided differences, depends very explicitly on the strict self-similarity of the measure and can therefore not be extended to the self-conformal setting.
 If $f=F_{\mu}$ is the distribution function of a Borel probability measure $\mu$, i.e.
-$F_{\mu}(x):=\mu\left((-\infty,x]\right)$ for $x\in\mathbb{R}$, then by [13, Lemma 5.1] we have
-(4) $\tilde{\alpha}_{F_{\mu}}(t_{0})=\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(t_{0},r)\right)}{\log r}.$
-For the remainder of this introduction, let $\mu$ a Gibbs measure associated with an arbitrary Hölder potential on the attractor of a conformal iterated function system (IFS) satisfying the open set condition - see the next section for precise definitions.
+$F_{\mu}(x):=\mu\left((-\infty,x]\right)$ for $x\in\mathbb{R}$, then by [9, Lemma 5.1] we have for each $t\in\mathbb{R}$,
+(2) $\tilde{\alpha}_{F_{\mu}}(t)=\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(t,r)\right)}{\log r}.$
+For the remainder of this introduction, let $\mu$ be a Gibbs measure associated with an arbitrary Hölder continuous potential on the attractor of a conformal iterated function system (IFS) satisfying the open set condition – see the next section for precise definitions.
 It is well known that the sets
@@
 satisfy the multifractal formalism; that is, $\dim_{H}\mathscr{E}(\alpha;\mu)=\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$, where $\beta^{*}(\alpha)$ is the Legendre transform of a suitable scaling function $\beta(q)$, defined in the next section.
+This result was first proved in [19] for conformal iterated functions, and in [21] for conformal repellers (see also [20] for a detailed account and related results).
 The second author and Sumi [9] showed recently that, in the case when all of the maps in the IFS are strictly increasing, the Hausdorff spectrum of the modified level sets
@@
 $\dim_{H}E_{F_{\mu}}(\alpha)=\beta^{*}(\alpha).$
-In a forthcoming paper, we will show that the situation is different for the packing spectrum.
+In a forthcoming paper, we will show that the situation is different for the packing spectrum, that is, the function $\alpha\mapsto\dim_{P}E_{f}(\alpha)$, where $\dim_{P}$ denotes the packing dimension.
+As an application of our main theorem, we determine the pointwise Hölder spectrum of conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps in Corollary 5.1.
+This result complements [9, Corollary 2.2], where the dimension spectrum of the conjugacy map is based on the simpler definition of the Hölder exponent $\tilde{\alpha}$.
+We note that it is straightforward to extend Corollary 5.1 to the framework of expanding circle endomorphisms.
+For related results on the non-differentiability of conjugacy maps in both frameworks we refer to [13].
+Remark 1.1.
+A closely related concept is that of $\alpha$-Hölder differentiability. For $\alpha>0$, say a function $f$ is $\alpha$-Hölder differentiable at a point $t_{0}$ if the limit
+$\lim_{t\to t_{0}}\frac{|f(t)-f(t_{0})|}{|t-t_{0}|^{\alpha}}$
+exists.
+The existence of the above limit (0, positive and finite, or $\infty$) gives information about the Hölder exponent $\tilde{\alpha}_{f}(t_{0})$ (but not, in general, about $\alpha_{f}(t_{0})$), and vice versa.
+Hölder differentiability for distribution functions of Gibbs measures was considered in [7, 14], with a later improvement in [24].
+Although some of the methods used in these papers (including the use of pressure functions and the bounded distortion principle) are similar to those used here, the principal questions are fundamentally different.
+This paper is organized as follows: In Section 2, we introduce the relevant notation and definitions.
+In Section 3, we state and prove our main result, namely the equality of $\alpha_{F}(t)$ and $\tilde{\alpha}_{F}(t)$ when $F$ is the distribution function of a Gibbs measure associated with a Hölder continuous potential $\psi$ on a self-conformal attractor.
+The only condition we impose on $\psi$ is that it is not cohomologous to the geometric potential $\varphi$ of the IFS; this is to avoid cases where $F$ is absolutely continuous and hence too “regular”.
+The technical heart of the proof is Proposition 3.7, which states that for each positive integer $k$ the limit
+$\lim_{y\to x}\frac{F(y)-F(x)}{(y-x)^{k}},$
+if it exists, cannot be nonzero and finite.
+As a consequence of the main result, we derive in Section 4 that the (Hausdorff) multifractal spectrum of such a distribution function is given by the multifractal formalism.
+In this section we show also that, at least in the self-similar case, the packing multifractal spectrum does not satisfy the multifractal formalism.
+Finally, in Section 5, we give an application of our main theorem to conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ maps.
 2. Notation and definitions
-Let $X\subset\mathbb{R}$ be a compact interval.
+Let $X\subset\mathbb{R}$ be a compact interval of positive length.
 A $\mathcal{C}^{1+\epsilon}$ conformal iterated function system on $X$ is given by an index set $I$ with $2\leq\#I<\infty$ and $\mathcal{C}^{1+\epsilon}$ conformal contracting maps $\Phi=(\phi_{i})_{i\in I}$, $\phi_{i}:X\rightarrow X$.
-More precisely, we mean that each $\phi_{i}$ extends to a $\mathcal{C}^{1+\epsilon}$ differentiable map on a neighborhood of $X$ which satisfies $0<\left|\phi_{i}^{\prime}(x)\right|<1$ for every $x\in X$ and for each $i\in I$.
-For ease of presentation, we assume that each of the maps $\phi_{i},i\in I$ is strictly increasing.
-However, this restriction is not essential.
+More precisely, we mean that, for each $i\in I$, the map $\phi_{i}$ extends to a $\mathcal{C}^{1+\epsilon}$ differentiable map on a neighborhood of $X$ which satisfies $0<\left|\phi_{i}^{\prime}(x)\right|<1$ for every $x\in X$.
 We refer to [16] for further details and basic properties of conformal iterated function systems.
-Let $\pi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ denote the coding map of $\Phi$ which is for $\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}$ given by
+Let $\pi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ denote the coding map of $\Phi$ which is for $\omega=(\omega_{1}\omega_{2}\dots)\in I^{\mathbb{N}}$ given by
 $\bigcap_{n\geq 1}\phi_{\omega_{1}}\circ\dots\circ\phi_{\omega_{n}}(X)=\left\{\pi(\omega)\right\}.$
@@
 We denote by $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ the left-shift on $I^{\mathbb{N}}$ which becomes a compact metric space endowed with the shift metric.
+A basis for the topology is given by the cylinder sets
+$w_{1}\dots w_{n}]:=\{\omega\in I^{\mathbb{N}}:\omega_{1}=w_{1},\dots,\omega_{n}=w_{n}\},\qquad w_{1}\dots w_{n}\in I^{n},\quad n\in\mathbb{N}.$
 Let $\varphi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be the geometric potential of $\Phi$, given by
-(5) $\varphi(\omega):=\log\left(\phi_{\omega_{1}}^{\prime}(\pi(\sigma(\omega)))\right),\quad\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}.$
-Since each $\phi_{i}$ is $\mathcal{C}^{1+\epsilon}$ differentiable, it follows that $\varphi$ is Hölder continuous. Let $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be Hölder continuous and denote by
+(3) $\varphi(\omega):=\log\left|\phi_{\omega_{1}}^{\prime}(\pi(\sigma(\omega)))\right|,\quad\omega=\omega_{1}\omega_{2}\dots\in I^{\mathbb{N}}.$
+Since each $\phi_{i}$ is $\mathcal{C}^{1+\epsilon}$ differentiable and the maps $\phi_{i}$ are uniformly contracting, it follows that $\varphi$ is Hölder continuous.
+Let $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be Hölder continuous and denote by
 $S_{n}\psi:=\sum_{k=0}^{n-1}\psi\circ\sigma^{k},\qquad n\geq 1$
-its ergodic sums.
+its ergodic sums. We will frequently make use of the fact that any Hölder continuous $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ satisfies the bounded distortion property
+(4) $\sup_{n\in\mathbb{N}}\sup_{w\in I^{n}}\sup_{\rho,\tau\in I^{\mathbb{N}}}\left|S_{n}\psi(w\rho)-S_{n}\psi(w\tau)\right|<\infty.$
 We say that a Borel probability measure $\mu_{\psi}$ on $I^{\mathbb{N}}$ is a Gibbs measure for $\psi$ if there exist two constants $C\geq 1$ and $P\in\mathbb{R}$ such that for each $n\in\mathbb{N}$ and for all $w\in I^{n}$ we have
-(6) $C^{-1}\leq\frac{\mu_{\psi}[w_{1},\dots,w_{n}]}{e^{S_{n}\psi(\overline{w})-nP}}\leq C.$
+(5) $C^{-1}\leq\frac{\mu_{\psi}[w_{1}\dots w_{n}]}{e^{S_{n}\psi(\overline{w})-nP}}\leq C.$
 Here $\overline{w}$ denotes the infinite periodic sequence with period block $w$.
@@
 $\mathcal{P}(\psi):=\lim_{k\to\infty}\frac{1}{k}\log\sum_{w\in I^{k}}\exp S_{k}\psi(\overline{w}).$
-We will frequently make use of the fact that any Hölder continuous $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ satisfies the bounded distortion property
-(7) $\sup_{n\in\mathbb{N}}\sup_{w\in I^{n}}\sup_{x,y\in I^{\mathbb{N}}}\left|S_{n}\psi(w\rho)-S_{n}\psi(w\tau)\right|<+\infty.$
+Two Hölder potentials $\psi$ and $\xi$ on $I^{\mathbb{N}}$ are said to be cohomologous if there is a bounded potential $\eta:I^{\mathbb{N}}\to\mathbb{R}$ such that $\psi-\xi=\eta-\eta\circ\sigma$.
+It is easy to see that, if $\psi$ and $\xi$ are cohomologous, then $\mathcal{P}(\psi)=\mathcal{P}(\xi)$.
 We remark that if $\mu_{\psi}$ is a Gibbs measure for $\psi$, then it is also a Gibbs measure for $\psi-\mathcal{P}(\psi)$.
@@
 (6) $\alpha_{+}:=\sup_{\omega\in I^{\mathbb{N}}}\limsup_{n\rightarrow\infty}\frac{S_{n}\psi(\omega)}{S_{n}\varphi(\omega)},\qquad\alpha_{-}:=\inf_{\omega\in I^{\mathbb{N}}}\liminf_{n\rightarrow\infty}\frac{S_{n}\psi(\omega)}{S_{n}\varphi(\omega)}.$
-It is well known that the multifractal spectrum $\mathcal{F}(\alpha)$ is complete [28], that is, $\mathcal{F}(\alpha)\neq\varnothing$ if and only if $\alpha\in\left[\alpha_{-},\alpha_{+}\right]$.
+It is well known that the multifractal spectrum is complete [22], that is, $\mathscr{E}(\alpha;\mu_{\psi})\neq\varnothing$ if and only if $\alpha\in\left[\alpha_{-},\alpha_{+}\right]$.
 Moreover, the following dichotomy is well known [21]: either we have $\alpha_{-}<\alpha_{+}$, or $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$.
 Note that in the latter case we have $\alpha_{-}=\alpha_{+}=\dim_{H}\Lambda$.
-Since $\varphi<0$, there is for each $q\in\mathbb{R}$ a unique $\beta(q)$ such that
+Since $\varphi<0$, there is for each $q\in\mathbb{R}$ a unique $\beta(q)\in\mathbb{R}$ such that
 (7) $\mathcal{P}(\beta(q)\varphi+q\psi)=0.$
@@
 Let $\Phi=(\phi_{i}:X\rightarrow X)_{i\in I}$ be a conformal iterated function system satisfying the open set condition.
+Let $\pi:I^{\mathbb{N}}\to\mathbb{R}$ denote the coding map of $\Phi$.
 Let $\varphi$ denote the geometric potential of $\Phi$ defined in (3) and let $\psi$ be any Hölder potential with $\mathcal{P}(\psi)=0$ such that $\psi$ is not cohomologous to $\varphi$.
@@
 Remark 3.2.
-The condition that $\psi$ is not cohomologous to $\varphi$ is needed to rule out some degenerate cases.
+The condition that $\psi$ is not cohomologous to $\varphi$ is needed to rule out cases where $F$ is too “smooth”.
 As a simple example, consider the IFS $\Phi=\{x/2,(x+1)/2\}$, whose attractor is the interval $[0,1]$.
@@
 On the other hand, it is easy to see that $\tilde{\alpha}_{F}(t)\equiv 1$ on $(0,1)$.
-More generally, we claim that for a Hölder continuous $\psi$ with $\mathcal{P}(\psi)=0$ we have that $\psi$ is cohomologous to $\varphi$ if and only if $\Lambda$ is an interval and some/any Gibbs measure $\mu_{\psi}$ for $\psi$ satisfies that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to the Lebesgue measure with a Radon-Nykodym density bounded away from zero and infinity.
-To prove this, first assume that $\psi$ is cohomologous to $\varphi$.
-Hence, $P(\varphi)=P(\psi)=0$ and thus, $\dim_{H}(\Lambda)=1$.
-Hence, by the well-known construction of conformal measures (see e.g.
-[19, Theorem 4.2.11]), $\Lambda$ has positive Lebesgue measure.
-Hence, by Lebesgue’s density theorem, we conclude that $X=\bigcup_{i\in I}\Phi_{i}(X)$ and thus, $\Lambda=X$ is an interval.
+More generally, we have the following characterization:
+Proposition 3.3.
+We have that $\psi$ is cohomologous to $\varphi$ if and only if $\Lambda$ is an interval and some (hence any) Gibbs measure $\mu_{\psi}$ for $\psi$ satisfies that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to the Lebesgue measure with a Radon-Nykodym density bounded away from zero and infinity.
+Proof.
+First assume that $\psi$ is cohomologous to $\varphi$.
+Hence, $P(\varphi)=P(\psi)=0$ and thus, by Bowen’s formula (see e.g.
+[17, Theorem 4.2.11]), $\Lambda$ has positive Lebesgue measure.
+We conclude that $X=\bigcup_{i\in I}\Phi_{i}(X)$ and therefore, $\Lambda=X$ is an interval.
 Indeed, if $X\setminus\bigcup_{i\in I}\Phi_{i}(X)$ was non-empty, it would contain an interval of positive length.
 By standard arguments, this would imply that $\Lambda$ contains no Lebesgue density points, which gives the desired contradiction.
-Consequently, the normalised Lebesgues measure on $\Lambda$ is a Gibbs measure for $\varphi$.
-Since $\psi$ is cohomologous to $\varphi$, $\mu_{\psi}$ is also a Gibbs measure for $\varphi$.
-It follows that $\mu_{\psi}$ is equivalent to the Lebesgue measure with a density bounded away from zero and infinity.
-Conversely, assume that $\Lambda$ is an interval and $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to Lebesgue with density bounded away from zero and infinity.
-Hence, Lebesgue measure is also Gibbs measure for $\psi$.
-Since Lebesgue measure is also a Gibbs measure for $\varphi$ and $P(\varphi)=P(\psi)=0$, it follows from [19, Theorem 2.2.7] that $\varphi$ is cohomologous to $\psi$.
-The key to proving Theorem 3.1 is to prove Proposition 3.4 below. The proof of that proposition uses the following lemma.
+Let $\lambda$ denote the normalised Lebesgue measure on $X$.
+Since $\Lambda=X$ and $P(\varphi)=0$, the unique Borel probability measure $\mu$ on $I^{\mathbb{N}}$ satisfying $\mu\circ\pi^{-1}=\lambda$ is a Gibbs measure for $\varphi$.
+Since $\psi$ is cohomologous to $\varphi$, it follows from [17, Theorem 2.2.7] that any Gibbs measure $\mu_{\psi}$ for $\psi$ is equivalent to $\mu$ with a density bounded away from zero and infinity.
+Hence, $\mu_{\psi}\circ\pi^{-1}$ is equivalent to $\lambda$ with a density bounded away from zero and infinity.
+Conversely, assume that $\Lambda$ is an interval.
+Then, $\Lambda=X$ and we have $P(\varphi)=0$ by Bowen’s formula.
+Further, the unique Borel probability measure $\mu$ on $I^{\mathbb{N}}$ defined by $\mu\circ\pi^{-1}=\lambda,$ is a Gibbs measure for $\varphi$.
+Assume that $\mu_{\psi}$ is a Gibbs measure for $\psi$ satisfying that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to $\lambda$ with density bounded away from zero and infinity.
+Then, $\mu$ is also a Gibbs measure for $\psi$.
+Since $P(\varphi)=P(\psi$), it follows from [17, Theorem 2.2.7] that $\varphi$ is cohomologous to $\psi$.
+∎
+The remainder of this section is devoted to the proof of Theorem 3.1.
+We will assume without loss of generality that the images of $X$ under the maps $\phi_{1},\dots,\phi_{m}$ are arranged from left to right; that is, $\inf\phi_{i+1}(X)\geq\sup\phi_{i}(X)$ for $i=1,\dots,m-1$, where $m:=|I|$.
+The first lemma is needed to deal with the codings of points that lie close to the common endpoint of adjacent cylinder intervals.
+The precise conclusion depends on the orientations of the first and last maps, $\phi_{1}$ and $\phi_{m}$, of the IFS.
+In this section, we write $\mathbf{i}|_{n}:=i_{1}\dots i_{n}$ for $\mathbf{i}=i_{1}\dots i_{p}\in I^{p}$ with $p\geq n$; and similarly, $\omega|_{n}:=\omega_{1}\dots\omega_{n}$ for $\omega=\omega_{1}\omega_{2}\dots\in I^{\mathbb{N}}$.
 Lemma 3.4.
-Suppose that $P(\psi)=0$. Suppose that there exists a finite set $A\subset\mathbb{R}$ and $\ell_{0}\geq 1$ such that for all $\ell\geq\ell_{0}$ and all $k_{1},\dots,k_{\ell}\in I$ we have
+Let $\mathbf{i},\mathbf{j}\in I^{p}$ with $\mathbf{i}\neq\mathbf{j}$, where $p\geq 3$.
+Let $n:=|\mathbf{i}\wedge\mathbf{j}|+1$ be the first coordinate at which $\mathbf{i}$ and $\mathbf{j}$ disagree.
+Assume $n\leq p-2$, and suppose the cylinder intervals $\pi([\mathbf{i}])$ and $\pi([\mathbf{j}])$ have a common endpoint.
+Then $|i_{n}-j_{n}|=1$ and one of the following holds:
+  1. (i)
+    $i_{n+1}\dots i_{p}\in\big\{1^{p-n},m^{p-n},1m^{p-n-1},m1^{p-n-1}\big\}$; or
+  2. (ii)
+    $i_{n+1}\dots i_{p}$ is a prefix of $(1m)^{\infty}$ and $i_{n}\neq m$; or
+  3. (iii)
+    $i_{n+1}\dots i_{p}$ is a prefix of $(m1)^{\infty}$ and $i_{n}\neq 1$.
+Furthermore, if $\mathbf{i}$ satisfies (ii) or (iii), then so does $\mathbf{j}$.
+Proof.
+Since the cylinder intervals $\pi([1]),\dots,\pi([m])$ are arranged from left to right, the cylinder intervals $\pi([\mathbf{i}|_{n-1}1]),\dots,\pi([\mathbf{i}|_{n-1}m])$ are arranged either from left to right or from right to left in $\pi([\mathbf{i}|_{n-1}])$ (depending on the orientation of $\phi_{i_{1}}\circ\dots\circ\phi_{i_{n-1}}$).
+Hence, $|i_{n}-j_{n}|=1$.
+Assume without loss of generality that $\pi([\mathbf{i}|_{n}])$ lies to the left of $\pi([\mathbf{j}|_{n}])$.
+Then for each $l=n+1,\dots,p$, $\pi([\mathbf{i}|_{l}])$ is the rightmost subcylinder of $\pi([\mathbf{i}|_{l-1}])$ and $\pi([\mathbf{j}|_{l}])$ is the leftmost subcylinder of $\pi([\mathbf{j}|_{l-1}])$.
+Thus, $i_{l},j_{l}\in\{1,m\}$ for $l=n+1,\dots,p$.
+If $\phi_{1}$ is increasing, then in $i_{n+1}\dots i_{p}$ any digit $1$ must be succeeded by another $1$.
+Similarly, if $\phi_{m}$ is increasing, any digit $m$ must be followed by another $m$.
+On the other hand, if $\phi_{1}$ is decreasing, any digit $1$ must be followed by a digit $m$; and likewise, if $\phi_{m}$ is decreasing, any digit $m$ must be followed by a digit $1$.
+This implies that $i_{n+1}\dots i_{p}$ is of one of the forms in (i) or else is a prefix of either $(1m)^{\infty}$ or $(m1)^{\infty}$.
+Suppose $i_{n+1}\dots i_{p}$ is a prefix of $(1m)^{\infty}$.
+Then $\phi_{1}$ and $\phi_{m}$ are both decreasing.
+If $i_{n}=m$, then necessarily $j_{n}=m-1$.
+Since $\pi([\mathbf{i}|_{n}])$ lies to the left of $\pi([\mathbf{j}|_{n}])$, this means that the map $\phi_{i_{1}}\circ\dots\circ\phi_{i_{n-1}}$ is decreasing.
+But $\phi_{m}$ is also decreasing, and so $\phi_{i_{1}}\circ\dots\circ\phi_{i_{n}}$ is increasing.
+Therefore, the rightmost subcylinder of $\pi([\mathbf{i}|_{n}])$ is $\pi([\mathbf{i}|_{n}m])$, contradicting that $i_{n+1}=1$.
+Hence, $i_{n}\neq m$.
+A similar argument shows that if $i_{n+1}\dots i_{n}$ is a prefix of $(m1)^{\infty}$, then $i_{n+1}\neq 1$.
+∎
+Lemma 3.5.
+Let $\ell\geq 2$ and let $\tau_{1}\dots\tau_{\ell}\in I^{\ell}$ be a word such that $\tau_{1}\neq\tau_{2}$.
+Then for any $\omega\in I^{\mathbb{N}}$ there are infinitely many integers $n$ such that the cylinder intervals $\pi([\omega|_{n+3\ell}])$ and $\pi([\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{3}])$ do not have a common endpoint.
+Proof.
+Let $\omega\in I^{\mathbb{N}}$, and consider two cases:
+Case 1.
+Suppose first that there exist infinitely many $n$ such that $\omega_{n}\neq\tau_{1}$.
+Take such an $n$.
+Suppose, by way of contradiction, that the cylinder intervals $\pi([\omega|_{n-1}\omega_{n}\dots\omega_{n+\ell+1}])$ and $\pi([\omega|_{n-1}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}])$ are adjacent.
+Then either (i), (ii) or (iii) of Lemma 3.4 holds for the word $\mathbf{i}:=\omega|_{n-1}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}$, with $p:=n+\ell+1$, so that $i_{n+1}\dots i_{p}=\tau_{2}\dots\tau_{\ell}\tau_{1}\tau_{2}$.
+Since $\tau_{2}\neq\tau_{1}$, we can rule out (i).
+If $\tau_{2}\dots\tau_{n}\tau_{1}\tau_{2}$ is a prefix of $(1m)^{\infty}$, then $\tau_{2}=1$ and so $\tau_{1}=m$, which means (ii) does not hold.
+By a symmetric argument, (iii) does not hold.
+Hence, $\pi([\omega|_{n+\ell+1}])$ and $\pi([\omega|_{n-1}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}])$ are not adjacent, so neither are $\pi([\omega|_{n-1+3\ell}])$ and $\pi([\omega|_{n-1}(\tau_{1}\dots\tau_{\ell})^{3}])$, since $\ell+2\leq 3\ell$.
+Case 2. Suppose $\omega$ ends in $\tau_{1}^{\infty}$. Then $\omega_{n-1}=\tau_{1}$ and $\omega_{n}\neq\tau_{2}$ for all large enough $n$. Fix such a large $n$. We have two further cases:
+Case 2a.
+Assume first that $\ell\geq 3$.
+Apply Lemma 3.4 with $p:=n+\ell+1$, $\mathbf{i}:=\omega|_{n-1}\tau_{2}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}=\omega|_{n-2}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}$ and $\mathbf{j}:=\omega|_{n+\ell+1}$, so $i_{n+1}\dots i_{p}=\tau_{3}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}$.
+We can immediately rule out condition (i) of Lemma 3.4, because $i_{n+1}=i_{p}$ and $i_{p-2}=\tau_{1}\neq\tau_{2}=i_{p-1}$.
+On the other hand, if $\tau_{3}\dots\tau_{n}\tau_{1}\tau_{2}\tau_{3}$ is a prefix of $(1m)^{\infty}$, then $\tau_{3}=1$ and hence $i_{n}=\tau_{2}=m$, so (ii) does not hold.
+Similarly, (iii) does not hold.
+Thus, the cylinder intervals $\pi([\mathbf{i}])$ and $\pi([\mathbf{j}])$ are not adjacent, so neither are $\pi([\omega|_{n-2+3\ell}])$ and $\pi([\omega|_{n-2}(\tau_{1}\dots\tau_{\ell})^{3}])$, since $\ell+3\leq 3\ell$.
+Case 2b. Finally, assume $\ell=2$. We then replace the word $\tau_{2}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}$ in Case 2a with $\tau_{2}\tau_{1}\tau_{2}\tau_{1}$, and argue similarly. ∎
+Lemma 3.6.
+Suppose that there exists a finite set $A\subset\mathbb{R}$ and $\ell_{0}\geq 1$ such that for all $\ell\geq\ell_{0}$ and all $k_{1},\dots,k_{\ell}\in I$ we have
 $\frac{S_{\ell}\psi\left(\overline{k_{1}\dots k_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{k_{1}\dots k_{\ell}}\right)}\in A.$
@@
 $\alpha_{-},\alpha_{+}]=\overline{\left\{\frac{S_{\ell}\psi\left(\overline{k_{1}\dots k_{\ell}}\right)}{S_{\ell}\varphi\left(\overline{k_{1}\dots k_{\ell}}\right)}:\ell\geq\ell_{0},k_{1},\dots,k_{\ell}\in I\right\}}.$
+Since $A$ is finite, we have $\alpha_{-}=\alpha_{+}$.
+By a well-known dichotomy for multifractal spectra [21], we have that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$.
+Clearly, this implies $A=\left\{\dim_{H}\Lambda\right\}$.
+∎
 Proposition 3.7.
-Assume that $\mathcal{P}(\psi)=0$, and $\psi$ is not cohomologous to $\varphi$. Then, for every $k\in\mathbb{N}$ and $x\in\mathbb{R}$ the limit
+For every $k\in\mathbb{N}$ and $x\in\mathbb{R}$ the limit
 $\lim_{y\rightarrow x}\frac{F(y)-F(x)}{(y-x)^{k}}$
 either does not exist, or it takes the value 0 or $\infty$.
-Proof of Proposition 3.4.
+Proof.
 The statement is trivial for even $k$, since $F$ is increasing. So below we will assume that $k$ is odd.
@@
 Note that the limit cannot be negative, since $F$ is increasing.
-Let $\omega\in I^{\mathbb{N}}$ such that $\pi(\omega)=x$. Let $s_{n},t_{n}]:=\pi\left(\left[\omega_{n}\right]\right)$. Then $s_{n}\leq x\leq t_{n}$, and
+Let $\omega\in I^{\mathbb{N}}$ such that $\pi(\omega)=x$. Let $s_{n},t_{n}]:=\pi\left(\left[\omega|_{n}\right]\right)$. Then $s_{n}\leq x\leq t_{n}$, and
 (9) $\begin{split}\frac{F(t_{n})-F(s_{n})}{(t_{n}-s_{n})^{k}}&=\left(\frac{t_{n}-x}{t_{n}-s_{n}}\right)^{k}\cdot\frac{F(t_{n})-F(x)}{(t_{n}-x)^{k}}\\ &\quad+\left(\frac{x-s_{n}}{t_{n}-s_{n}}\right)^{k}\cdot\frac{F(s_{n})-F(x)}{(s_{n}-x)^{k}}.\end{split}$
@@
 Indeed such $\tau_{1}\dots\tau_{\ell}$ always exist for otherwise, Lemma 3.6 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=k=1$, contradicting our hypothesis.
+Replacing $\tau_{1}\dots\tau_{\ell}$ by a cyclical permutation if necessary, we may assume that $\tau_{2}\neq\tau_{1}$.
 Taking a subsequence if necessary, we may assume that $\lim r_{n}=r\in[0,1]$. Then by (9), we have
@@
 $r_{n}^{N}:=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}.$
-Recall that the maps $\phi_{i},i\in I$ are all increasing, so if $u$ and $v$ are two words of the same length and $u\prec v$ in the lexicographical order, then $\pi([u])$ lies to the left of $\pi([v])$.
-We claim that there is an infinite subset $\mathcal{N}\subset\mathbb{N}$ such that for each $n\in\mathcal{N}$, the point $x$ and the interval $s_{n}^{N},t_{n}^{N}]$ are separated by at least a full cylinder interval of level at most $n+\ell+1$.
-To see this, let $j<\ell$ be the smallest integer such that $\tau_{j+1}\neq\tau_{1}$.
-Such $j$ exists since $\tau_{1}\dots\tau_{\ell}$ contains at least two different letters.
-Write $\tau_{1}\dots\tau_{\ell}=\tau_{1}^{j}\tau_{j+1}\dots\tau_{\ell}$, and consider two cases:
-Case 1.
-If $\omega$ ends in $\tau_{1}^{\infty}$, then for every sufficiently large $n\in\mathbb{N}$ the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}^{j+1}\tau_{j+1}]\big{)}$ separates $x$ from the interval $\pi\big{(}[\omega|_{n}\tau_{1}\dots\tau_{j+1}]\big{)}$, and hence from $s_{n}^{N},t_{n}^{N}]$.
-Case 2. Otherwise, there are infinitely many $n$ such that $\omega_{n+1}\neq\tau_{1}$. Take such an $n$. If $\omega_{n+1}<\tau_{1}$, then
-$\omega_{n+1}\dots\omega_{n+\ell+1}\prec\tau_{1}1^{\ell}\prec\tau_{1}\dots\tau_{\ell}\tau_{1},$
-so the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}1^{\ell}]\big{)}$ lies between $x$ and $s_{n}^{N},t_{n}^{N}]$.
-A similar argument, with the inequalities reversed, applies when $\omega_{n+1}>\tau_{1}$.
-In either case, the claim follows.
-By replacing $\mathcal{N}$ by an infinite subset if necessary, we may assume that either $t_{n}^{N}>x$ for all $n\in\mathcal{N}$, or $t_{n}^{N}<x$ for all $n\in\mathcal{N}$.
+By Lemma 3.5 there is an infinite subset $\mathcal{N}\subset\mathbb{N}$ such that for each $n\in\mathcal{N}$, the point $x$ and the interval $s_{n}^{N},t_{n}^{N}]$ are separated by a full cylinder interval of level $n+3\ell$.
+Replacing $\mathcal{N}$ by an infinite subset if necessary, we may assume that either $t_{n}^{N}>x$ for all $n\in\mathcal{N}$, or $t_{n}^{N}<x$ for all $n\in\mathcal{N}$.
 Without loss of generality, we assume the former.
@@
 Let $r_{\min}>0$ be a number such that $|\phi_{i}^{\prime}(x)|\geq r_{\min}$ for all $i\in I$ and $x\in X$. Then
-(13) $\mathrm{diam}\pi([uv])\geq\mathrm{diam}\pi([u])\cdot r_{\min}^{|v|}$
+(11) $\mathrm{diam}(\pi([uv]))\geq\mathrm{diam}(\pi([u]))\cdot r_{\min}^{|v|}$
 for any two finite words $u$ and $v$.
-For $n\in\mathcal{N}$, let $u_{n},v_{n}]$ be a cylinder interval of length $n+\ell+1$ separating $x$ and $s_{n}^{N},t_{n}^{N}]$. Then by (13),
-$|t_{n}^{N}-x|\geq v_{n}-u_{n}\geq\mathrm{diam}\pi([\omega|_{n}])\cdot r_{\min}^{\ell+1}.$
+For $n\in\mathcal{N}$, let $u_{n},v_{n}]$ be a cylinder interval of length $n+3\ell$ separating $x$ and $s_{n}^{N},t_{n}^{N}]$. Then by (11),
+$|t_{n}^{N}-x|\geq v_{n}-u_{n}\geq\mathrm{diam}(\pi([\omega|_{n}]))\cdot r_{\min}^{3\ell}.$
 Hence, there is a constant $C=C(\ell)>1$, depending only on $\ell$ but not on $n$ or $N$, such that
@@
 Now (12) implies
-(15) $r_{n}^{N}=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}\asymp_{\ell}\frac{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)}{\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}\asymp_{\ell}e^{-NS_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
+(13) $r_{n}^{N}=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}\asymp_{\ell}\frac{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)}{\mathrm{diam}(\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}\asymp_{\ell}e^{-NS_{\ell}\left(\varphi\right))(\overline{\tau_{1}\dots\tau_{\ell}})}.$
 Hence, by passing to a subsequence, using a diagonal argument, we may assume that $r_{n}^{N}$ converges to some number $r^{N}\in\mathbb{R}$ for every $N\geq 1$.
@@
 (14) $\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}=\left((r^{N})^{k}+(1-r^{N})^{k}\right)L\geq 2^{1-k}L.$
-On the other hand, by the Gibbs property (6) of $\mu_{\psi}$ and the bounded distortion property (12) we have
-$\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}$ $=\frac{\mu_{\psi}\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)\right)^{k}}$
-$\asymp_{\ell}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}\right]\right)\right)^{k}}\cdot e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}$
+On the other hand, by the Gibbs property (5) of $\mu_{\psi}$, the bounded distortion property (4) and (10) we have
+$\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}$ $=\frac{\mu_{\psi}\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}{\left(\mathrm{diam}(\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right])\right)\right)^{k}}$
+$\asymp_{\ell}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\left(\mathrm{diam}(\pi\left(\left[\omega|_{n}\right])\right)\right)^{k}}\cdot e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}$
 $\asymp_{\ell}e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
@@
 Suppose now that $k\geq 3$. We have shown that
-$2^{1-k}L\leq\big{(}(r^{N})^{k}+(1-r^{N})^{k}\big{)}L=\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}\asymp e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
+$2^{1-k}L\leq\big((r^{N})^{k}+(1-r^{N})^{k}\big)L=\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}\asymp_{\ell}e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
 Since $S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)\neq 0$ it follows that $S_{\ell}(\psi-k\varphi)\left(\overline{\tau_{1}\dots\tau_{\ell}}\right)>0$, and hence, $r^{N}\rightarrow\infty$ as $N\rightarrow\infty$.
@@
 4. The pointwise Hölder spectrum of $F_{\mu}$
-Recall from the Introduction that $E_{F}(\alpha)=\{t\in\mathbb{R}:\alpha_{F}(t)=\alpha\}$. As a consequence of Theorem 3.1, we obtain the following result.
+Recall from the Introduction that $E_{F}(\alpha)=\{t\in\mathbb{R}:\alpha_{F}(t)=\alpha\}$.
+As a consequence of Theorem 3.1, we obtain the following result, in which $\alpha_{-}$ and $\alpha_{+}$ are given by (6), and $\beta^{*}(\alpha)$ is defined by (8).
 Corollary 4.1.
@@
 But we also have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\leq\beta^{*}(\alpha)$.
-This follows directly from a careful examination of the proof that $\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$; see, for instance, [21, p.
+This follows directly from a careful examination of the proof that $\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$; see, for instance, [19, p.
 501].
@@
 In a forthcoming paper, we will extend the packing spectrum (17) to the case of distribution functions of general Gibbs measures.
+5. Conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ maps
+We apply our results to conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps, extending [9, Corollary 2.3] in two ways.
+Firstly, we also analyze the pointwise Hölder exponent $\alpha$ of the conjugacy map, whereas in [9] only $\tilde{\alpha}$ was considered.
+Secondly, we allow interval maps with orientation-reversing branches.
+Let $f$ be an expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval map with $s\geq 2$ full branches, i.e., there exist closed intervals $J_{1},\dots,J_{s}\subset[0,1]$ with non-empty, pairwise disjoint interiors such that $f|_{J_{i}}$ has a $\mathcal{C}^{1+\epsilon}$ extension to a neighbourhood of $J_{i}$ satisfying $|f^{\prime}|_{J_{i}}|>1$ and $f(J_{i})=[0,1]$, for $1\leq i\leq s$.
+We will always assume that $\sup J_{i}\leq\inf J_{j}$ if $i<j$.
+The repeller of $f$ is denoted by $\Lambda$ and the restriction $f|_{\Lambda}:\Lambda\rightarrow\Lambda$ is semi conjugate to the shift $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ with alphabet $I:=\left\{1,\dots,s\right\}$.
+The semi conjugacy is given by the coding map $\pi_{f}:I^{\mathbb{N}}\rightarrow\Lambda$ of the conformal iterated function systems $\Phi_{f}$, which is given by the contracting inverse branches $(f|_{J_{i}})^{-1}:[0,1]\rightarrow[0,1]$, $1\leq i\leq s$.
+Note that each branch is either orientation-preserving, or orientation-reversing.
+We denote by $\varphi_{f}:I^{\mathbb{N}}\rightarrow\mathbb{R}$ the geometric potential of $\Phi_{f}$ given by $\varphi_{f}:=-\log|f^{\prime}\circ\pi_{f}|$.
+We say that two piecewise $\mathcal{C}^{1+\epsilon}$ interval maps $f$ and $g$ have the same signature if they share the same alphabet $I$ and satisfy
+$\{i\in I\mid f_{i}\text{ is orientation-preserving }\}=\{i\in I\mid g_{i}\text{ is orientation-preserving }\}.$
+If $f$ and $g$ have the same signature and $\Lambda_{g}=[0,1]$, then $\pi_{g}(\pi_{f}^{-1}\{x\})$ is a singleton for each $x\in\Lambda_{f}$. We can thus define the conjugacy map
+$\Theta:\Lambda_{f}\rightarrow[0,1],\quad\{\Theta(x)\}=\pi_{g}(\pi_{f}^{-1}\{x\}).$
+Slightly abusing notation, we write $\Theta=\pi_{g}\circ\pi_{f}^{-1}$. Note that we have $\Theta\circ f=g\circ\Theta$, which implies for each $x\in\Lambda_{f}$,
+(18) $\pi_{g}^{-1}\left((-\infty,\Theta(x)]\right)=\pi_{f}^{-1}\left((-\infty,x]\right).$
+The conjugacy $\Theta$ has a unique continuous extension to a function on $\mathbb{R}$ which is locally constant on $\mathbb{R}\setminus\Lambda_{f}$.
+We denote this function again by $\Theta$ and refer to it as the conjugacy map between $f$ and $g$.
+Note that (18) continues to hold for the conjugacy map between $f$ and $g$.
+Since $\Lambda_{g}=[0,1]$ we have that the Lebesgue measure $\lambda$ on $[0,1]$ satisfies $\lambda=\mu_{\varphi_{g}}\circ\pi_{g}^{-1}$ where $\mu_{\varphi_{g}}$ is a Gibbs measure for $\varphi_{g}$.
+Hence, by (18), the conjugacy map $\Theta$ between $f$ and $g$ coincides with the distribution function of $\mu_{\varphi_{g}}\circ\pi_{f}^{-1}$.
+Our main theorem thus implies the following.
+Corollary 5.1.
+Let $f$ and $g$ be two expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps with full branches and with the same signature.
+Suppose that $\varphi_{f}$ is not cohomologous to $\varphi_{g}$.
+Then, for the conjugacy map $\Theta:\Lambda_{f}\rightarrow[0,1]$ between between $f$ and $g$ we have
+$\dim_{H}\left\{t\in\Lambda\mid\alpha_{\Theta}(t)=\alpha\right\}=\dim_{H}\left\{t\in\Lambda\mid\tilde{\alpha}_{\Theta}(t)=\alpha\right\}=\beta^{*}(\alpha),\quad\alpha\in\left[\alpha_{-},\alpha_{+}\right],$
+where $\beta:\mathbb{R}\rightarrow\mathbb{R}$ is defined implicitly by $\mathcal{P}(\beta(q)\varphi_{f}+q\varphi_{g})=0$, and $\alpha_{-}<\alpha_{+}$.
+Remark 5.2.
+Let $f$ and $g$ be two expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps with full branches.
+Suppose in addition that all branches of $f$ and $g$ are monotonically increasing and that $\Lambda_{f}=\Lambda_{g}=[0,1]$.
+Then, it follows from [13, Proof of Theorem 1.2] that the conjugacy map $\Theta:[0,1]\rightarrow[0,1]$ between $f$ and $g$ is a $\mathcal{C}^{1+\epsilon}$-diffeomorphism if and only if $\varphi_{f}$ is cohomologous to $\varphi_{g}$.
+Example 5.3.
+Let
+$f(x):=\begin{cases}3x&\mbox{if $0\leq x\leq 1/3$},\\ 3x-2&\mbox{if $2/3\leq x\leq 1$},\end{cases}\qquad g(x):=\begin{cases}2x&\mbox{if $0\leq x<1/2$},\\ 2x-1&\mbox{if $1/2\leq x\leq 1$}.\end{cases}$
+Then the repeller $\Lambda$ of $f$ is the ternary Cantor set and the conjugacy $\Theta$ between $f$ and $g$ is the ternary Cantor function.
+Suppose we invert the second branch of both $f$ and $g$; that is, we consider instead the functions
+$\tilde{f}(x):=\begin{cases}3x&\mbox{if $0\leq x\leq 1/3$},\\ 4-3x&\mbox{if $2/3\leq x\leq 1$},\end{cases}\qquad\tilde{g}(x):=\begin{cases}2x&\mbox{if $0\leq x<1/2$},\\ 3-2x&\mbox{if $1/2\leq x\leq 1$}.\end{cases}$
+Note that $\tilde{f}$ and $\tilde{g}$ again have the same signature.
+The repeller of $\tilde{f}$ is still the ternary Cantor set, and the conjugacy $\tilde{\Theta}$ between $\tilde{f}$ and $\tilde{g}$ is still the ternary Cantor function.
+In other words, $\tilde{\Theta}:=\pi_{\tilde{g}}\circ\pi_{\tilde{f}}^{-1}=\Theta$.
+This follows because, if the coding of $x\in\Lambda$ under $f$ is $\omega:=(\omega_{n})_{n}:=\pi_{f}^{-1}(x)$, then the coding of $x$ under $\tilde{f}$ is $\tilde{\omega}=(\tilde{\omega}_{n})_{n}=\pi_{\tilde{f}}^{-1}(x)$, where $\tilde{\omega}_{1}=\omega_{1}$ and successively, for $n=2,3,\dots$,
+$\tilde{\omega}_{n}:=\begin{cases}\omega_{n}&\mbox{if $\#\{i<n:\tilde{\omega}_{i}=2\}$ is even},\\ 3-\omega_{n}&\mbox{otherwise},\end{cases}$
+[…]
+Example 5.4.
+More generally, let $m,n\in\mathbb{N}$ with $m,n\geq 3$. Let
+$f(x):=\begin{cases}mx&\mbox{if $0\leq x\leq 1/m$},\\ nx-(n-1)&\mbox{if $(n-1)/n\leq x\leq 1$},\end{cases}$
+and let $g(x)$ again be as in Example 5.3.
+If $m=n$, the situation is much as it was in the last example.
+But if $m\neq n$, the conjugacy $\Theta$ changes when we reverse the second branch of both $f$ and $g$, because the repeller $\Lambda$ changes.
+Nonetheless, the new conjugacy $\tilde{\Theta}$ has the same Hölder spectrum as $\Theta$.
+The range of the spectrum is
+$\alpha_{-},\alpha_{+}]=\left[\min\left\{\frac{\log 2}{\log m},\frac{\log 2}{\log n}\right\},\max\left\{\frac{\log 2}{\log m},\frac{\log 2}{\log n}\right\}\right],$
+and the spectrum is given by Corollary 5.1.
+Acknowledgments
+P.A. was partially supported by Simons Foundation grant #709869. J.J. was partially supported by the JSPS KAKENHI 24K06777.